Analysis of 20.0 g of a compound containing only calcium and bromine indicates that 4.00 g of calcium are present. What is the empirical formula of the compound formed?

A step by step solution if possible, thank you!
1
Expert's answer
2020-11-09T13:57:58-0500

Ca = 4g

Br = (20-4)g = 16g


Firstly we divide by their molar masses

Ca = 4/40 = 0.1

Br = 16/80 = 0.2


Then we divide by the lowest number

Ca = 0.1/0.1 = 1

Br = 0.2/0.1 = 2


"\therefore" the empirical formula of the compound formed is "CaBr_2"